find the measure of two complementary angles if one is 10 degrees more than three times the other
Answer by Cromlix(4381)

Other angle = x
Larger angle = 10 + 3x
Complementary angles = 90 degrees
So, x + (10 + 3x) = 90
x + 10 + 3x = 90
4x = 90 - 10
4x = 80
x = 20
Other angle = 20 degrees
Larger angle = 70 degrees.
Hope this helps.
:-)